OpenStudy (anonymous):

Write an equation for the line in point/slope form and slope/intercept form that has the given condition. 1a. Slope = ½ and passes through the origin.

OpenStudy (anonymous):

@iGreen

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the origin is the point \((0,0)\)

OpenStudy (misty1212):

so just \[y=\frac{1}{2}x\] is all

OpenStudy (anonymous):

Slope Intercept Form : y=mx+b Use the point slope formula which is y-y1 = m(x-x1) A. y-y1 = m(x-x1) m=1/2 y1=0 x1= 0 So, its y-0 = 1/2 (x-0) It will be y = 1/2x y = 1/2x

OpenStudy (anonymous):

Haha hi misty1212, thank you for helping you guys. Could you do two more?

OpenStudy (anonymous):

sure